OpenStudy (aundiewear):

Simplify ((cosθ)/(1+sinθ) - (cosθ)/(sinθ -1))^-1 a) cosθ/2 b) 2secθ c) 2sinθ d) cscθ/2

OpenStudy (kittiwitti1):

\[\frac{\cos{θ}}{1+\sin{θ}}-(\frac{\cos{θ}}{\sin{θ}-1})^{-1}?\]

OpenStudy (kittiwitti1):

\[(\frac{a}{b})^{-1}=\frac{b}{a}\]

OpenStudy (aundiewear):

the ^-1 is for the whole equation

OpenStudy (aundiewear):

oh so I flip both parts?

OpenStudy (kittiwitti1):

oh well then you'll have to solve by greatest common multiple I think. make it one fraction

OpenStudy (kittiwitti1):

(1+sinθ)(sinθ-1) denominator,

OpenStudy (kittiwitti1):

sorry I've got homework of my own so I'm not sure if I can stay here

OpenStudy (aundiewear):

oh it's okay thank you for your help (:

OpenStudy (kittiwitti1):

np. do you get what I'm saying though? make a common denominator ?

OpenStudy (aundiewear):

yes i do

OpenStudy (kittiwitti1):

alright, do you think you can do it from here? I might be able to help but I'm not certain

OpenStudy (aundiewear):

you just multiply the denominator to both sides right?

OpenStudy (kittiwitti1):

well the left side multiplies by (1+sinθ/1+sinθ) and the right fraction multiplies by (sinθ-1/sinθ-1)

OpenStudy (aundiewear):

yeah

OpenStudy (kittiwitti1):

then you find any identities in the resulting fraction, and simplify

OpenStudy (kittiwitti1):

that's how I learned to do it ☺

OpenStudy (aundiewear):

im actually not sure what the answer should be. I got "a" as the answer

OpenStudy (kittiwitti1):

oh. I'm sorry I don't have enough time to help you u_u @Directrix ?

OpenStudy (anonymous):

\[\frac{1}{\frac{\cos (\theta )}{\sin (\theta )+1}-\frac{\cos (\theta )}{\sin (\theta )-1}}=-\frac{1}{\frac{2 \cos (\theta )}{(\sin (\theta )-1) (\sin (\theta )+1)}}=-\frac{1}{\frac{2 \cos (\theta )}{\sin ^2(\theta )-1}}=-\frac{1}{-\frac{2 \cos (\theta )}{\cos ^2(\theta )}}=\frac{\cos (\theta )}{2} \]